Write two equations in two variables and solve.
A motorboat traveling with the current went 72 miles in 4 hours. Going against the current, it took 6 hours to travel the same distance. Find the rate of the boat in calm water and the rate of the current.

Answer by ankor@dixie-net.com(22740)

A motorboat traveling with the current went 72 miles in 4 hours.
Going against the current, it took 6 hours to travel the same distance.
Find the rate of the boat in calm water and the rate of the current.
:
let s = the rate of the boat in calm water
let c = the rate of the current
:
Write a distance equation for each way
4(s+c) = 72
6(s-6) = 72
Simplify both equations, divide the 1st by 4 and the 2nd by 6, you have
s + c = 18
s - c = 12
-------------adding eliminates c, find s
2s = 30
s = 15 mph in calm water
:
I'll let you find the current rate, check solution is both original equations